OpenStudy (anonymous):

combining like terms simplify each expression: 15n -19n = -4n

OpenStudy (lilg132):

n(15-19) = -4n?

OpenStudy (lilg132):

n(15-19) = n(-4)

OpenStudy (anonymous):

simplify each expression

OpenStudy (anonymous):

15n -19n = -4n -4n = -4n this is an identity

OpenStudy (anonymous):

jimmyrep am i right?

OpenStudy (anonymous):

what you have written is correct - yes n(15-19) = n(-4) further simplification gives what i wrote

OpenStudy (anonymous):

ok thank you

OpenStudy (lilg132):

but that simplification doesnt make sense -4n = -4n thats just like saying 0 = 0 i dont know maybe im wrong

OpenStudy (anonymous):

I am combining like terms and simplifying each expression

OpenStudy (anonymous):

well 0 is = 0 15n -19n = -4n is an identity not an equation because it is true for all real values of n

myininaya (myininaya):

i think his problem says simplify 15n-19n and he is saying it simplifies to -4n

myininaya (myininaya):

which is correct

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes that is what i am saying

OpenStudy (lilg132):

looks like i wasn't wrong but jimmyrep simplified it better then me :p